# Ameliorating transient noise bursts in gravitational-wave searches for intermediate-mass black holes

### v1

@@ V.2 Training with known data set of O3a
Once we have selected $t_{w}$, and after several fine-tuning experiments, we can train and validate our model with the known data set of O3a. As we mentioned in Section IV.2, we use 9-fold cross-validation to enhance the performance of the algorithm. Thus, in Fig. 9 we show, for training and validation, the mean accuracy (top panel) and loss (bottom panel) of the 9-fold cross-validation as a function of the epochs, where the shaded region represents ±$\pm 3$ standard deviations. Note that the loss is plotted in logarithmic scale so we can appreciate the difference between training and validation loss is minimal, which implies that the model is not overfitting. Another relevant point is that both mean accuracy and loss seem to have sharp peaks around epochs 900 and 1700, which could imply that the learning is unstable. However, this is a known effect of the adaptive learning rate that we described in Section IV.3. The described behavior is also present in L1 and V1, whose results can be found in Appendix A.2.
@@ V.3 Diving into the known data: machine learning performance
Employing GPU Tesla V100 with a memory of 161616\,16Gb allowed us to train our model in $\approx 17.40\,$h for H1 data, $\approx 15.55\,$h for L1 data, $\approx 31.20\,$h for V1 data. Such a large amount of time is mainly due to the 9-fold cross-validation procedure. Nonetheless, once the models are trained we can predict a single input in $2.9\times 10^{-6}\,$s. As we also want to test the generalization power of the model between O3a and O3b known data set, we present their confusion matrix for H1 in Fig. 10 (see results for L1 and V1 in Appendix A.3). While in the y-axis we represent the ground truth, in the x-axis we represent the predictions of the model. Furthermore, the percentages in the rows sum $100\%$, and the elements in the diagonal were successfully classified.
@@
To assess the degree of misclasification during O3b with respect to the SNR distribution we compute the TPR as a function of the average SNR $\bar{\rho}$ (Fig. 11a). For this aim, we arrange the Injections in SNR bins and compute their TPR setting the decision threshold to be $P_{inj}^{*}=0.9$. Interestingly, we can observe that the TPR is high for low $\bar{\rho}$, showing dip at $\bar{\rho}\sim 8$ of TPR0.85similar-toabsent0.85\,$\sim$ 0.85∼ 0.85 for H1 , and TPR0.7similar-toabsent0.7\,$\sim$ 0.7∼ 0.7 for L1 and V1. Moreover, the TPR becomes unstable at $\bar{\rho}>15$.
To explain this behavior we need to observe the SNR $\rho$ distributions, computed by Omicron Robinet et al. (2020), of the different classes for the known data set of O3b (Fig. 11b). Firstly, it is important to note that Gravity Spy classifies glitches with Omicron SNR $>7.5$ Glanzer et al. (2023). In Fig. 11b we show the distribution of $\rho$, the resulting SNR from Omicron, where we can still observe that most glitch populations are centred around $\rho\sim 8$, which explains the dip in Fig. 11a. Our method understands that most classes of problematic glitches, such as Fast_Scattering, Tomte and Low_Frequency_Burst are around this value, so it does not have trouble distinguishing Injections at lower $\rho$. Note that at $\rho<5$ the model simply differentiates between Whistle, Fast_Scattering and Injection classes. Furthermore, as $\rho$ increases, the population of Injections decreases, so due to the low statistic at $\bar{\rho}>15$ the TPR seems unstable. We note that a different SNR distribution for the Injection class might lead to a different behavior of the TPR as a function of the SNR. We will leave this exploration to future work.
@@
A possible cause for this lack of generalization between O3a and O3b can be the fact that the interferometers are evolving systems, so the glitches produced at the beginning of the observing run can be different from the ones produced at the end (see Fig. 4 of Abbott et al. (2023)). Furthermore, we remind the reader that while the Injections class is the actual ground truth, there is some bias in the definition of the glitch classes (e.g. Fig. 8). Another reason is the limitation of the model itself, as we are performing a multi-classification task with information on 6 variables (see Eq. 9). To lower the number of FPs, i.e. glitches that are incorrectly classified as Injections, in this single-detector task we can use time coincidence between detectors (see Section IV.4 for a definition), since it is less common for glitches to appear in several interferometers at the same time.
@@ V.4 Diving into the known data: significance of $P_{inj}$ statistic
Figure 13: ROC curve of double and triple detector time (H1L1V1) for Eq. 14 (dashed lines) and Eq. 15. Here, TPR is the rate of correctly classified GW injections from the zero lag, while FPR is the rate of incorrectly classified time slides background tracks.
Up until now, we performed the standard statistical tests in the field of ML with $P_{inj}$ as our probability. Nonetheless, a ML based detection algorithm requires a proper understanding of the significance of its statistics considering the accidental background. With the time-shifted background, we intend to find extreme values of $\bar{P}_{inj}$ (see Eq. 13), but due to the probabilistic nature of ML algorithms, our statistic saturates at 1 as we can see in Fig. 12. Since we are interested in the region with the highest probability we redefine our statistics as
$-\log{(1-\bar{P}_{inj})+\epsilon},$ (14)
where $\bar{\cdot}$ represents again the harmonic mean (see Eq. 13 for a definition), and $\epsilon=10^{-20}$ is added to prevent undefined behavior when $\bar{P}_{inj}\approx 1$, ensuring numerical stability, firstly proposed in Trovato et al. (2024). The main advantage of using the harmonic mean in this context is that it places more emphasis on smaller values, which helps to down-rank candidates with lower statistics. To assess the performance of this statistic we compute the ROC curve, but with a different definition of TP and FP. TP are coincident GW injections of O3a present in the zero lag that has been identified by GstLAL, i.e. they produce a track. On the other hand, FP are the coincident tracks of the time shifts of O3. With this definition we can construct the ROC curve stepping on different values of our statistic $-\log{(1-\bar{P}_{inj})+\epsilon}$. Note that better performing ROC would be the ones that maximize the area under the curve (AUC). It is relevant to note that, as we are working with triple time if one detector does not observe the signal, then its contribution $P^{-1}_{inj,k}=0$ (see Eq. 13).
@@
In Fig. 13 we present the ROC curve of $-\log{(1-\bar{P}_{inj})+\epsilon}$ (dashed lines) in triple detector time (H1L1V1) for different detector combinations. We can observe that H1L1 and H1V1 have a better performance than L1V1. This implies that the L1V1 combination has more background tracks, possibly due to a larger glitch population in both detectors for different reasons. On the other hand, H1L1V1 has a better performance than double-coincidences since we only have two background tracks. We summarize their AUC in Table 3.
@@
where again $\bar{\cdot}$ represents the harmonic mean. We present in Fig. 13 (solid lines) its ROC curve in triple detector time (H1L1V1). We can observe that all double coincidences have improved greatly with respect to the statistic $-\log{(1-\bar{P}_{inj})+\epsilon}$. Moreover, in triple coincidence, the AUC is 9877.87, while before it was 9838.81 (see Table 3).
[…]
Figure 14: Inverse false alarm rate (iFAR) as a function of the ranking statistic in triple time (H1L1V1) for different detector combinations. The solid lines show the time-shifted background with real GW signal tracks, while dashed lines show the time-shifted background without them.
A principal characteristic of a GW ranking statistic is that it increases monotonically as the probability of finding a GW signal increases. Because of this, real GW signal tracks would lay on the far right of the iFAR distribution, increasing the false alarm. To investigate whether our ranking statistic behaves in this way we remove all tracks that correspond to real GW signals, yielding the dashed line distributions. While there is a noticeable discrepancy in the H1L1 distribution, this issue does not occur with H1V1 and L1V1. This may be due to the lower sensitivity of V1 or a limitation of our statistics.
@@
Figure 15: TPR, i.e., the rate of correctly classified GW injections, at iFAR = 0.1 (equivalent to one per 10 years) for different detector combinations in triple time (H1L1V1). Error bars represent the standard error at 3 standard deviations.
To further understand the performance of our statistic, we measure the recovery of GW injections from O3a. To compute the TPR, we fix iFAR at 0.1, or equivalently, one per 10 years, using the results in Fig. 14. In Fig. 15 we present TPR as a function of the injected SNR (Fig. 15a) and as a function of the luminosity distance $D_{L}$ in Mpc (Fig. 15b).
@@ VII Conclusion
    Construction of accidental background: We constructed the time shift background to assess the significance of our ranking statistic (see Section V.4) and tested its performance on simulated IMBH signals. With an iFAR $=0.1\,$yr, we have a TPR $\gtrsim 0.8$ at SNR $\sim 5$ for all detector combination. Regarding the luminosity distance, H1L1 and H1L1V1 have a TPR $\gtrsim 0.8$, with a drop in performance at $D_{L}\gtrsim 1.2\,$Mpc.
@@
In future work, we will explore a different ML method that can process varying length inputs, as the time information might be relevant to enhance the performance of the model. For this aim, several ML algorithms could be employed, such as recurrent neural network Staudemeyer and Rothstein Morris (2019) or even transformers with an attention mechanism Vaswani and toehrs (2017). With such a model we could perform a fair comparison with state-of-the-art pipelines to quantify the potential improvements in IMBH searches that this technology may offer. Furthermore, and as we mentioned before, this methodology is flexible and simply relies on matched filtering computation. We could therefore extend it to other CBC signals and state-of-the-art matched-filtering algorithms.

### v2

@@ V.2 Training with known data set of O3a
Once we have selected $t_{w}$, and after several fine-tuning experiments, we can train and validate our model with the known data set of O3a. As we mentioned in Section IV.2, we use 9-fold cross-validation to enhance the performance of the algorithm. Thus, in Fig. 9 we show, for training and validation, the mean accuracy (top panel) and loss (bottom panel) of the 9-fold cross-validation as a function of the epochs, where the shaded region represents $\pm 3$ standard deviations. Note that the loss is plotted in logarithmic scale so we can appreciate the difference between training and validation loss is minimal, which implies that the model is not overfitting. Another relevant point is that both mean accuracy and loss seem to have sharp peaks around epochs 900 and 1700, which could imply that the learning is unstable. However, this is a known effect of the adaptive learning rate that we described in Section IV.3. The described behavior is also present in L1 and V1, whose results can be found in Appendix A.2.
@@ V.3 Diving into the known data: machine learning performance
Employing GPU Tesla V100 with a memory of 1616\,Gb allowed us to train our model in $\approx 17.40\,$h for H1 data, $\approx 15.55\,$h for L1 data, $\approx 31.20\,$h for V1 data. Such a large amount of time is mainly due to the 9-fold cross-validation procedure. Nonetheless, once the models are trained we can predict a single input in $2.9\times 10^{-6}\,$s. As we also want to test the generalization power of the model between O3a and O3b known data set, we present their confusion matrix for H1 in Fig. 10 (see results for L1 and V1 in Appendix A.3). While in the y-axis we represent the ground truth, in the x-axis we represent the predictions of the model. Furthermore, the percentages in the rows sum $100\%$, and the elements in the diagonal were successfully classified.
@@
[…]
To explain this behavior we need to observe the SNR $\rho$ distributions, computed by Omicron Robinet:2020lbf, of the different classes for the known data set of O3b (Fig. 11b). Firstly, it is important to note that Gravity Spy classifies glitches with Omicron SNR >7.5>7.5 Glanzer:2022avx. In Fig. 11b we show the distribution of $\rho$, the resulting SNR from Omicron, where we can still observe that most glitch populations are centred around $\rho\sim 8$, which explains the dip in Fig. 11a. Our method understands that most classes of problematic glitches, such as Fast_Scattering, Tomte and Low_Frequency_Burst are around this value, so it does not have trouble distinguishing Injections at lower $\rho$. Note that at $\rho<5$ the model simply differentiates between Whistle, Fast_Scattering and Injection classes. Furthermore, as $\rho$ increases, the population of Injections decreases, so due to the low statistic at $\bar{\rho}>15$ the TPR seems unstable. We note that a different SNR distribution for the Injection class might lead to a different behavior of the TPR as a function of the SNR. We will leave this exploration to future work.
@@
A possible cause for this lack of generalization between O3a and O3b can be the fact that the interferometers are evolving systems, so the glitches produced at the beginning of the observing run can be different from the ones produced at the end (see Fig. 4 of LIGOScientific:2021djp). Furthermore, we remind the reader that while the Injections class is the actual ground truth, there is some bias in the definition of the glitch classes (e.g. Fig. 8). Another reason is the limitation of the model itself, as we are performing a multi-classification task with information on 6 variables (see Eq. 9). To lower the number of FPs, i.e. glitches that are incorrectly classified as Injections, in this single-detector task we can use time coincidence between detectors (see Section IV.4 for a definition), since it is less common for glitches to appear in several interferometers at the same time.
@@ V.4 Diving into the known data: significance of $P_{inj}$ statistic
Figure 13: ROC curve of double and triple detector time (H1L1V1) for Eq. 14 (dashed), Eq. 15 (solid) and Eq. 16 (dotted). Here, TPR is the rate of correctly classified GW injections from the zero lag, while FPR is the rate of incorrectly classified time slides background tracks.
Up until now, we performed the standard statistical tests in the field of ML with $P_{inj}$ as our classification probability. Nonetheless, a ML based detection algorithm requires a proper understanding of the significance of its statistics considering the accidental background. With the time-shifted background, we intend to find extreme values of $\bar{P}_{inj}$ (see Eq. 13), but due to the probabilistic nature of ML algorithms, our statistic saturates at 1 as we can see in Fig. 12. Since we are interested in the region with the highest probability we redefine our statistics as
$\log{(1-\bar{P}_{inj})+\epsilon},$ (14)
where $\bar{\cdot}$ represents again the harmonic mean (see Eq. 13 for a definition), and $\epsilon=10^{-20}$ is added to prevent undefined behavior when $\bar{P}_{inj}\approx 1$, ensuring numerical stability, firstly proposed in Trovato:2023bby. The main advantage of using the harmonic mean in this context is that it places more emphasis on smaller values, which helps to down-rank candidates with lower statistics. To assess the performance of this statistic we compute the ROC curve, but with a different definition of TP and FP. TP are coincident GW injections of O3a present in the zero lag that has been identified by GstLAL, i.e. they produce a track. On the other hand, FP are the coincident tracks of the time shifts of O3. With this definition we can construct the ROC curve stepping on different values of our statistic $\log{(1-\bar{P}_{inj})+\epsilon}$. Note that better performing ROC would be the ones that maximize the area under the curve (AUC). It is relevant to note that, as we are working with triple time if one detector does not observe the signal, then its contribution $P^{-1}_{inj,k}=0$ (see Eq. 13).
@@
$\overline{\rho_{\text{MF}}}$
H1L1 8795.28
H1V1 8548.83
L1V1 8686.17
H1L1V1 0.00
In Fig. 13 we present the ROC curve of $\log{(1-\bar{P}_{inj})+\epsilon}$ (dashed lines) in triple detector time (H1L1V1) for different detector combinations. We can observe that H1L1 has a better performance than H1V1 and L1V1. This could be due to a known lower sensitivity of V1 detector. However, L1V1 is worse performing than H1V1. A possible explanation for this is that L1 contains more glitches due to its higher sensitivity, confussing our ML model.
On the other hand, H1L1V1 has a better performance than double-coincidences since we only have two background tracks. We summarize their AUC in Table 3.
@@
where again $\bar{\cdot}$ represents the harmonic mean. We present in Fig. 13 (solid lines) its ROC curve in triple detector time (H1L1V1). We can observe that all double coincidences have improved greatly with respect to the statistic $\log{(1-\bar{P}_{inj})+\epsilon}$. Moreover, in triple coincidence, the AUC is 9877.87, while before it was 9838.81 (see Table 3).
To further understand the performance of our ML-enhanced statistic compared to standard matched filtering, we compare it to an approximate detection statistic defined in Magee:2023muf and 2024arXiv241015513M as
$\rho_{\text{MF}}=\frac{\rho}{[\frac{1}{2}(1+\max{(1,\xi^{2})}^{3})]^{1/5}}$ (16)
for a single detector. As before, we use the harmonic mean to combine the different interferometers, yielding $\overline{\rho_{\text{MF}}}$ statistic. The resulting ROC curves are shown in Fig. 13 (dotted). We can see that $\rho_{\text{MF}}$ outperforms $\log{(1-\bar{P}_{inj})+\epsilon}$ at low FPR, but it has a worse performance than $\overline{P_{inj}/(\xi^{2}/\text{SNR}^{2})}$. This can also be observed in Table 3, where all the values of $\overline{\rho_{\text{MF}}}$ are lower than $\overline{P_{inj}/(\xi^{2}/\text{SNR}^{2})}$. It is interesting to note that $\overline{\rho_{\text{MF}}}$ for three detectors is zero, which might be caused due to the lack of sufficient background statistic for this combination of detectors.
False alarm rate (FAR) is typically used to assess the performance of a GW search. This is measured by time-shifting the detectors’ data relative to each other, as described in Sec. III.4. Figure 14 shows the FAR as a function of our ranking statistic $\overline{P_{inj}/(\xi^{2}/\text{SNR}^{2})}$ for different detector combinations in H1L1V1 time. Note that H1L1V1 is in the bottom left corner as it has two samples.
Figure 14: False alarm rate (FAR) as a function of the ranking statistic in triple time (H1L1V1) for different detector combinations. The solid lines show the time-shifted background with real GW signal tracks, while dashed lines show the time-shifted background without them.
A principal characteristic of a GW ranking statistic is that it increases monotonically as the probability of finding a GW signal increases. Because of this, real GW signal tracks would lay on the far right of the FAR distribution. To investigate whether our ranking statistic behaves in this way we remove all tracks that correspond to real GW signals, yielding the dashed line distributions. While there is a noticeable discrepancy in the H1L1 distribution, this issue does not occur with H1V1 and L1V1. This may be due to the lower sensitivity of V1 or a limitation of our statistics.
@@
Figure 15: TPR, i.e., the rate of correctly classified GW injections, at FAR = 0.1 (equivalent to one per 10 years) for different detector combinations in triple time (H1L1V1). Error bars represent the standard error at 3 standard deviations.
To further understand the performance of our statistic, we measure the recovery of GW injections from O3a. To compute the TPR, we fix FAR at 0.1, or equivalently, one per 10 years, using the results in Fig. 14. In Fig. 15 we present TPR as a function of the injected SNR (Fig. 15a) and as a function of the luminosity distance $D_{L}$ in Mpc (Fig. 15b).
@@ VII Conclusion
    Construction of accidental background: We constructed the time shift background to assess the significance of our ranking statistic (see Section V.4) and tested its performance on simulated IMBH signals. With an FAR =0.1=0.1\,yr, we have a TPR $\gtrsim 0.8$ at SNR $\sim 5$ for all detector combination. Regarding the luminosity distance, H1L1 and H1L1V1 have a TPR $\gtrsim 0.8$, with a drop in performance at $D_{L}\gtrsim 1.2\,$Mpc.
@@
In future work, we will explore a different ML method that can process varying length inputs, as the time information might be relevant to enhance the performance of the model. For this aim, several ML algorithms could be employed, such as recurrent neural network lstm or even transformers with an attention mechanism attention. With such a model we could perform a fair comparison with state-of-the-art pipelines to quantify the potential improvements in IMBH searches that this technology may offer. Furthermore, and as we mentioned before, this methodology is flexible and simply relies on matched filtering computation. We could therefore extend it to other CBC signals and state-of-the-art matched-filtering algorithms.